no beam no physics last night - start setup now

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.gov)
Date: Thu Feb 27 2003 - 16:25:28 EST

  • Next message: Apache: "Shift report 20030228 16:00-24:00"
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    
    Another update from the sched. physicist
    
    Flemming
    
    
    Sent: Thursday, February 27, 2003 3:55 PM
    Subject: no beam no physics last night - start setup now
    
    
    | Hi there,
    |
    | (you are right, it's not a good sign to get another mail so quickly after the last one .. ;)).
    |
    | finally! - there's beam in RHIC again and we are working on injection. About 9:00 pm everybody was swept last night and
    we were ready for beam. This is
    | when problems started with cooling leaks in the AtR which kept haunting us until about 3:00 pm today :(. PP setup
    continued during the maintenance day in
    | booster and AGS. BLIP was off all day yesterday and during the night.
    |
    | RHIC RF group did do all landau cavity tests which could be done without beam during the day and is now ready to do a 6
    bunch test ramp. After that we'll
    | work on a 56 bunch production ramp with high bunch intensities. Landau cavity work will be continued tomorrow during the
    day.
    |
    | Good luck!
    |
    |     Angelika
    |
    |
    |
    


    This archive was generated by hypermail 2.1.5 : Thu Feb 27 2003 - 16:25:41 EST